17 May 2010

While we wait for the next plane tomorrow

Perfectly primed skiway
Jakob has taken this beautiful shot of our skiway.

Today was the first day of real drilling. While still in the education mode, the new drillers are learning fast. Now 15 m core is waiting in the buffer.

Processors are unpacking equipment and paperwork and are eager to begin. However, we need the last pieces of equipment which are scheduled on the plane tomorrow. This equipment can be in place in one day, so we expect processing to begin on Wednesday or Thursday.

On the plane tomorrow is also fresh vegetables and other food stuff. We look forward to it. We cross our fingers, that weather holds and that the air crew finds the skiway in good order.

What we have done today:
1. Drilling and logging.
2. Drilling school.
3. Setting up equipment in new warm laboratory.
4. Fixing problem with drillers centrifuge and drillers electronics.
5. Seismic station maintained and operational. It has been leveled and centered.
6. Cleaning up in the cargo line and removing snow drifts from camp.
7. Japanese pit-study continues.
8. Setting up water vapour sampling site.
9. Setting up CFA laboratory.

Ad.1: Drillers report:
Production in the drill trench is ready. Two shifts have been formed consisting of Trevor, Martin and Vasileios in the mornings and Romain, Adrian and Jack in the evenings with the exceptionally strong support of Jakob and Steff handling the big issues as they arise.
After 24 filter runs picking up 226 kg of excess chips from the 585 runs from last season we are drilling in a clean hole. Most of the chips were picked up in the upper 800 meters of the borehole meaning the chips either float or remain suspended in the liquid, leaving the bottom relatively clean while drilling. In 6 drilling runs beginning Saturday afternoon we are picking up and average of 21.7 kg of chips per 3.5 meters of ice drilled.

So far drilling has been smooth. The reasons for two short runs were easily solved. One was due to an anti-torque rotation as we conservatively set the blades a little too weak for the first run. The second due to an electronic failure in the surface modem, that has been repaired.
The drill configuration has been set to exactly match the configuration that we successfully ended with last year.

Today, 4 runs for 11.00 meters. Driller's cable depth 1757.8.

Weather: blue sky, -24°C to - 14°C, 8 knots from SE. Visibility: Unrestricted.

FL, J.P. Steffensen

← Previous entry   Next entry →